# Semantic-ICP: Iterative Closest Point for Non-rigid Multi-Organ Point Cloud Registration

### v2

@@ 3.2 Compared Baselines
We compare our method with established rigid and non-rigid point-matching-based point cloud registration methods. The baseline methods include the original ICP [6], Global-ICP (GO-ICP) [13], Fast-robust-ICP (FR-ICP) [14], non-rigid ICP (NR-ICP) [10], Fast-RNRR [11] and AMM-NRR [12]. We used the default parameters provided in the baselines’ open-source codes. To evaluate the point cloud surface similarity, we use the 95th percentile of the Hausdorff Distance (HD95) and the mean surface distance (MSD) to determine the surface differences between the source and target point clouds.
@@ 3.4 Evaluation on Datasets
We report the quantitative evaluation using HD95 and MSD across three datasets, as shown in Tables 1 to 6. Compared to other baselines, our method SemICP achieves the lowest HD95 and MSD for each anatomical structure, as well as the lowest average across labels, on all three datasets.
TORS: The quantitative results are shown in Table 1 and 2. AMM-NRR and Fast-RNRR completed registration on only 7 pairs of 24 data pairs, with numerical issues affecting the remaining data. One Fast-RNRR result was highly inaccurate and is treated as a failed registration, not included in the final analysis. One unique challenge associated with this dataset is that the US provides only partial observation of relevant anatomical structures, especially the larynx and bony structures, resulting in differing point cloud sizes between US and CT. Additionally, visibility varies across anatomical structures. Our results show that other registration methods have larger HD95 and MSD when registering all anatomical structures, especially the larynx and bony structures. Without semantic information, other non-rigid registration algorithms under-penalize the larynx/bone registration error. Mismatching different structures also occurs in other baselines without semantic information, as shown in the qualitative evaluation in Fig. 2. Our method improves the delineation of anatomical structures and leads to significantly lower HD95 and MSD across all labels, greatly improving registration accuracy.
@@
Initial ICP GO-ICP FR-ICP SemICP-rigid nr-ICP AMM-NRR Fast-RNRR SemICP
Carotid $5.47\pm 2.68$ $9.63\pm 5.18$ $5.68\pm 3.47$ $7.84\pm 4.95$ $4.59\pm 2.52$ $9.12\pm 5.78$ $17.4\pm 0.40$ $3.01\pm 1.58$ $\mathbf{0.61\pm 0.17}$
Larnyx/bone $10.21\pm 9.27$ $8.07\pm 8.44$ $20\pm 9.04 7.7$$78\pm 8.92$ $7.18\pm 8.62$ $5.57\pm 9.43$ $4.10\pm 3.24$ $5.92\pm 5.08$ $\mathbf{1.12\pm 2.18}$
Vein $10.85\pm 4.29$ $7.24\pm 5.22$ $5.47\pm 3.01$ $7.17\pm 5.69$ $5.38\pm 2.62$ $6.38\pm 6.08$ $3.08\pm 0.72$ $6.53\pm 5.12$ $\mathbf{0.68\pm 0.37}$
Average $8.84$ $8.31$ $6.12$ $7.60$ $5.72$ $7.02$ $2.97$ $5.15$ $\mathbf{0.80}$
Table 2: The average MSD on the TORS dataset. The metrics are reported in mean (std) in mmmm and the best performance is highlighted in bold.
Initial ICP GO-ICP FR-ICP SemICP-rigid nr-ICP AMM-NRR Fast-RNRR SemICP
Carotid $2.19\pm 0.92$ $4.02\pm 2.61$ $2.07\pm 1.08$ $2.87\pm 1.83$ $1.86\pm 0.92$ $3.55\pm 2.91$ $0.71\pm 0.10$ $1.25\pm 0.66$ $\mathbf{0.30\pm 0.05}$
Larnyx/bone $4.72\pm 4.26$ $3.44\pm 5.29$ $3.23\pm 5.41$ $3.40\pm 5.68$ $3.07\pm 4.78$ $2.47\pm 5.59$ $1.70\pm 2.16$ $2.74\pm 3.34$ $\mathbf{0.34\pm 0.29}$
Vein $5.25\pm 2.52$ $3.37\pm 2.64$ $2.39\pm 1.46$ $2.93\pm 2.31$ $2.37\pm 1.28$ $2.65\pm 3.01$ $0.93\pm 0.12$ $3.09\pm 2.44$ $\mathbf{0.32\pm 0.09}$
Average $4.05$ $3.61$ $2.56$ $3.07$ $2.43$ $2.89$ $1.11$ $2.36$ $\mathbf{0.32}$
Figure 2: Visualization of the registration results on one example US-CT pair in TORS dataset (posterior-anterior view).
AbdominalMRCT: Table 3 and 4 show and HD95 and MSD on this dataset. Results for AMM-NRR and Fast-RNRR are reported on only 7 MR-CT pairs due to a numerical problem encountered with one point cloud pair. The challenge of this dataset is that the organs undergo significant deformation, so all rigid registration methods have a relatively large surface distance as shown in Fig. 3. Our method achieves the lowest surface distance for all organs compared to baseline methods.
@@
Initial ICP GO-ICP FR-ICP SemICP-rigid nr-ICP AMM-NRR Fast-RNRR SemICP
@@
Spleen $37.12\pm 17.25$ $16.89\pm 5.78$ $16.06\pm 6.08$ $24.29\pm 11.60$ $16.73\pm 5.58$ $3.55\pm 3.39$ $6.87\pm 4.58$ $10.45\pm 13.59$ $\mathbf{1.06\pm 0.05}$
[…]
Kidney-L $33.15\pm 14.37$ $20.80\pm 9.24$ $1.13\pm 9.65 22$$2.70\pm 10.06$ $21.17\pm 9.80$ $5.66\pm 3.74$ $17.71\pm 17.01$ $12.68\pm 12.85$ $\mathbf{1.32\pm 0.21}$
Average $35.69$ $17.98$ $17.81$ $21.98$ $17.90$ $4.30$ $13.21$ $14.66$ $\mathbf{1.17}$
Table 4: The average MSD on the Learn2reg AbdominalMRCT. The metrics are reported in mean (std) in mmmm and the best performance is highlighted in bold.
Initial ICP GO-ICP FR-ICP SemICP-rigid nr-ICP AMM-NRR Fast-RNRR SemICP
Liver $18.22\pm 7.00$ $8.31\pm 3.39$ $8.14\pm 3.20$ $9.07\pm 3.82$ $8.15\pm 3.26$ $1.73\pm 1.41$ $2.00\pm 0.69$ $6.86\pm 8.61$ $\mathbf{0.64\pm 0.05}$
Spleen $18.22\pm 11.84$ $7.36\pm 2.51$ $7.26\pm 2.72$ $9.65\pm 5.52$ $7.62\pm 2.62$ $1.49\pm 1.13$ $2.33\pm 1.35$ $4.71\pm 6.83$ $\mathbf{0.63\pm 0.03}$
Kidney-R $17.05\pm 12.99$ $6.83\pm 3.59$ $6.82\pm 3.14$ $5.64\pm 2.66$ $6.48\pm 3.21$ $1.71\pm 1.59$ $10.88\pm 17.96$ $9.26\pm 12.39$ $\mathbf{0.69\pm 0.06}$
Kidney-L $14.32\pm 6.04$ $8.86\pm 4.05$ $8.98\pm 4.05$ $69\pm 4.46 8.8$$83\pm 4.01$ $2.25\pm 1.59$ $4.84\pm 4.16$ $5.01\pm 5.80$ $\mathbf{0.73\pm 0.05}$
Average $16.95$ $7.84$ $7.80$ $8.26$ $7.77$ $1.80$ $5.01$ $6.46$ $\mathbf{0.67}$
Figure 3: Visualization of the registration results on one example CT-MR pair in AbdominalMRCT dataset. The first row is the anterior-posterior view, and the second row is the posterior-anterior view.
HippocampusMR: HD95 and MSD results are reported in Table 5 and Table 6, respectively. This dataset includes only the head and tail of the hippocampus, which are relatively small structures with limited deformation compared to abdominal organs. While other non-rigid methods also achieve sub-millimeter HD95 and MSD, our method SemICP yields the lowest distance errors. AMM-NRR and Fast-RNRR completed registration on 58 pairs, with two pairs excluded due to numerical issues while building the deformation graph and sampling the influencing node sets.
Initial ICP GO-ICP FR-ICP SemICP-rigid nr-ICP AMM-NRR Fast-RNRR SemICP
Head $3.68\pm 1.46$ $2.44\pm 1.05$ $2.46\pm 1.12$ $2.84\pm 1.43$ $2.34\pm 0.90$ $0.85\pm 0.76$ $0.85\pm 0.76$ $2.61\pm 1.60$ $\mathbf{0.28\pm 0.03}$
Tail $3.68\pm 1.23$ $23\pm 0.97 2.2$$21\pm 1.06$ $2.46\pm 1.29$ $1.97\pm 0.79$ $0.82\pm 0.75$ $0.82\pm 0.75$ $2.12\pm 1.45$ $\mathbf{0.28\pm 0.02}$
Average $3.68$ $2.33$ $2.33$ $2.64$ $2.16$ $0.84$ $0.84$ $2.36$ $\mathbf{0.28}$
Table 6: The average MSD on the Learn2reg hippocampusMR. The metrics are reported in mean (std) in mmmm.
Initial ICP GO-ICP FR-ICP SemICP-rigid nr-ICP AMM-NRR Fast-RNRR SemICP
Head $1.73\pm 0.73$ $1.04\pm 0.36$ $1.04\pm 0.37$ $1.09\pm 0.47$ $1.03\pm 0.34$ $0.40\pm 0.21$ $0.40\pm 0.21$ $0.95\pm 0.50$ $\mathbf{0.12\pm 0.02}$
Tail $1.60\pm 0.48$ $0.94\pm 0.29$ $0.93\pm 0.31$ $0.91\pm 0.31$ $0.88\pm 0.26$ $0.40\pm 0.23$ $0.40\pm 0.23$ $0.76\pm 0.31$ $\mathbf{0.12\pm 0.01}$
Average $1.67$ $0.99$ $0.98$ $1.00$ $0.95$ $0.40$ $0.40$ $0.85$ $\mathbf{0.12}$
Figure 4: Visualization of the registration results on one example MR-MR pair in HippocampusMR dataset (superior-inferior view).

### v3

@@ 3.2 Compared Baselines
We compare our method with different types of established point cloud registration methods. The first type of methods are rigid and non-rigid point-matching-based point cloud registration, which are the most similar methods to our method. The baseline methods include: the original ICP (Besl and McKay (1992)), Global-ICP (GO-ICP) (Yang et al. (2015)), Fast-robust-ICP (FR-ICP) (Zhang et al. (2022)), non-rigid ICP (NR-ICP) (Amberg et al. (2007)), Fast-RNRR (Yao et al. (2020)), AMM-NRR (Yao et al. (2023)). The second type of method is probabilistic-based registration, and we choose to use Coherent Point Drift (CPD) (Gatti and Khallaghi (2022)) to provide a GMM-based point cloud registration baseline. To compare with learning-based methods, we choose to use the two Deep Closest Point (DCP) models described in (Wang and Solomon (2019)) trained on ModelNet40 (Wu et al. (2015)). DCPv1 uses a Dynamic Graph CNN as the embedding model to learn the feature representation of the points, and performs matching on the features directly, while DCPv2 introduces a transformer block for the point matching. We used the default parameters provided in the open-source codes of the baselines. For ICP, GO-ICP, and CPD, we downsample the point cloud to 1000 points to accelerate their run time. For DCP, we follow the paper’s implementation, downsampling the point cloud to 1024 points, and normalizing them to a uniform ball.
To evaluate the point cloud surface similarity, we use the 95th percentile of the Hausdorff Distance (HD95) and the mean surface distance (MSD) to determine the surface differences between the source and target point clouds.
@@ 3.4 Evaluation on Datasets
We report the quantitative evaluation using HD95 and MSD across three datasets, as shown in Tables 1 to 3. Compared to other baselines, our method SemICP achieves the lowest HD95 and MSD for each anatomical structure, as well as the lowest average across labels, on all three datasets.
TORS: The quantitative results are shown in Table 1. AMM-NRR and Fast-RNRR completed registration on only 7 pairs of 24 data pairs, with numerical issues affecting the remaining data. One Fast-RNRR result was highly inaccurate and is treated as a failed registration, not included in the final analysis. One unique challenge associated with this dataset is that the US provides only partial observation of relevant anatomical structures, especially the larynx and bony structures, resulting in differing point cloud sizes between US and CT. Additionally, visibility varies across anatomical structures. Our results show that other registration methods have larger HD95 and MSD when registering all anatomical structures, especially the larynx and bony structures. Without semantic information, other non-rigid registration algorithms under-penalize the larynx/bone registration error. Mismatching different structures also occurs in other baselines without semantic information, as shown in the qualitative evaluation in Fig. 2. Our method improves the delineation of anatomical structures and leads to significantly lower HD95 and MSD across all labels, greatly improving registration accuracy. It is worth noted that deep learning-based methods, DCPv1 and DCPv2, increases the surface matching error, showing that these models can not be applied in an out-of-box manner for TORS dataset. Our method also performs registration faster than other non-rigid registration methods, with an average computation time as 34.51 seconds, showing the potential to be utilized in image guidance systems to perform registration.
@@
HD95 MSD Runtime
Carotid Larnyx/bone Vein Average Carotid Larnyx/bone Vein Average
Initial $5.47\pm 2.68$ $10.21\pm 9.27$ $10.85\pm 4.29$ $8.84$ $2.19\pm 0.92$ $4.72\pm 4.26$ $5.25\pm 2.52$ $4.05$
ICP $6.22\pm 2.99$ $7.07\pm 10.11$ $5.41\pm 2.99$ $6.24$ $2.47\pm 1.25$ $2.95\pm 5.59$ $2.32\pm 1.39$ $2.58$ $\mathbf{0.04\pm 0.01}$
GO-ICP $6.22\pm 3.56$ $5.53\pm 8.21$ $\pm 2.35 5.605.60$ $2.44\pm 1.29$ $1.73\pm 2.26$ $2.22\pm 1.04$ $2.13$ $6.85\pm 1.37$
[…]
CPD $10.70\pm 2.86$ $9.48\pm 7.43$ $8.87\pm 6.65$ $9.68$ $4.47\pm 1.17$ $3.67\pm 3.22$ $3.86\pm 3.72$ $4.00$ $2.78\pm 0.85$
DCPv1 $27.32\pm 8.30$ $15.53\pm 9.01$ $26.58\pm 13.70$ $23.14$ $12.97\pm 4.84$ $9.49\pm 6.91$ $15.59\pm 10.76$ $12.36$ $0.14\pm 0.40$
DCPv2 $25.63\pm 7.60$ $15.52\pm 9.49$ $5.62\pm 13.53 22$.2522.25 $12.02\pm 4.45$ $9.45\pm 7.73$ $4.22\pm 11.08 11$.8911.89 $0.15\pm 0.40$
nr-ICP $8.27\pm 5.25$ $5.90\pm 9.41$ $5.98\pm 5.44$ $6.72$ $3.38\pm 2.52$ $2.54\pm 5.36$ $2.66\pm 2.73$ $2.86$ $165.49\pm 243.38$
AMM-NRR $17.4\pm 0.40$ $4.10\pm 3.24$ $3.08\pm 0.72$ $2.97$ $0.71\pm 0.10$ $1.70\pm 2.16$ $0.93\pm 0.12$ $1.11$ $894.98\pm 188.62$
Fast-RNRR $3.01\pm 1.58$ $5.92\pm 5.08$ $6.53\pm 5.12$ $5.15$ $1.25\pm 0.66$ $2.74\pm 3.34$ $3.09\pm 2.44$ $2.36$ $2408.30\pm 3755.40$
SemICP $\mathbf{0.63\pm 0.18}$ $\mathbf{1.17\pm 2.33}$ $\mathbf{0.69\pm 0.40}$ $\mathbf{0.83}$ $\mathbf{0.31\pm 0.05}$ $\mathbf{0.36\pm 0.35}$ $\mathbf{0.32\pm 0.10}$ $\mathbf{0.33}$ $34.51\pm 16.96$
Figure 2: Visualization of the registration results on one example US-CT pair in TORS dataset (posterior-anterior view). The medial-lateral view is shown in supplemental material Fig. S1.
AbdominalMRCT: Table 2 show and HD95 and MSD on this dataset. Results for AMM-NRR and Fast-RNRR are reported on only 7 MR-CT pairs due to a numerical problem encountered with one point cloud pair. The challenge of this dataset is that the organs undergo significant deformation, so all rigid registration methods have a relatively large surface distance as shown in Fig. 3. DCP models achieve the lowest registration accuracy due to the domain mismatch, and it is worth noting that it will be hard to train a model since we have only 8 pairs of data. Our method achieves the lowest surface distance for all organs compared to baseline methods, and the registration time is approximately 33 seconds, which is acceptable for intraoperative systems.
@@
HD95 MSD Runtime
Liver Spleen Kidney-R Kidney-L Average Liver Spleen Kidney-R Kidney-L Average
Initial $38.07\pm 14.60$ $37.12\pm 17.25$ $4.41\pm 20.42 33$$3.15\pm 14.37$ $35.69$ $18.22\pm 7.00$ $18.22\pm 11.84$ $17.05\pm 12.99$ $14.32\pm 6.04$ $16.95$
ICP $22.02\pm 7.41$ $15.00\pm 4.34$ $13.12\pm 4.91$ $17.25\pm 6.95$ $16.85$ $9.41\pm 3.60$ $6.67\pm 2.38$ $5.82\pm 2.38$ $7.08\pm 3.14$ $7.24$ $\mathbf{0.04\pm 0.01}$
GO-ICP $21.95\pm 7.84$ $14.80\pm 4.45$ $12.41\pm 5.03$ $\pm 6.67 16.4716.47$ $9.19\pm 3.86$ $6.54\pm 2.27$ $5.47\pm 3.47$ $6.99\pm 2.95$ $7.05$ $16.39\pm 20.29$
FR-ICP $26.35\pm 10.78$ $24.29\pm 11.60$ $14.59\pm 7.86$ $22.70\pm 10.06$ $21.98$ $9.07\pm 3.82$ $9.65\pm 5.52$ $5.64\pm 2.66$ $8.69\pm 4.46$ $8.26$ $155.87\pm 60.02$
CPD $20.53\pm 9.73$ $24.19\pm 18.81$ $1.93\pm 4.96 11$$1.00\pm 4.91$ $16.91$ $8.53\pm 3.28$ $11.06\pm 9.33$ $5.09\pm 1.64$ $4.32\pm 1.29$ $7.25$ $2.82\pm 0.52$
DCPv1 $59.76\pm 14.24$ $88.37\pm 34.20$ $71.61\pm 15.10$ $66.82\pm 30.57$ $71.64$ $21.74\pm 3.94$ $55.94\pm 29.14$ $9.57\pm 12.73 33$$3.34\pm 18.90$ $37.65$ $0.83\pm 1.86$
DCPv2 $62.70\pm 26.93$ $126.96\pm 60.90$ $98.70\pm 48.30$ $81.23\pm 41.01$ $92.40$ $25.97\pm 10.92$ $91.49\pm 61.39$ $65.54\pm 39.96$ $51.57\pm 34.15$ $58.65$ $0.83\pm 1.86$
nr-ICP $5.80\pm 6.20$ $5.14\pm 6.33$ $6.56\pm 8.28$ $10.15\pm 11.01$ $6.91$ $2.45\pm 2.72$ $2.15\pm 2.47$ $2.71\pm 3.52$ $4.04\pm 4.62$ $2.84$ $543.08\pm 336.46$
AMM-NRR $6.05\pm 2.25$ $6.87\pm 4.58$ $22.20\pm 27.03$ $17.71\pm 17.01$ $13.21$ $2.00\pm 0.69$ $2.33\pm 1.35$ $10.88\pm 17.96$ $4.84\pm 4.16$ $5.01$ $664.61\pm 211.11$
Fast-RNRR $16.95\pm 17.57$ $10.45\pm 13.59$ $18.55\pm 20.74$ $12.68\pm 12.85$ $14.66$ $6.86\pm 8.61$ $4.71\pm 6.83$ $9.26\pm 12.39$ $5.01\pm 5.80$ $6.46$ $367.09\pm 168.03$
SemICP $\mathbf{1.15\pm 0.16}$ $\mathbf{1.05\pm 0.05}$ $\mathbf{1.23\pm 0.27}$ $\mathbf{1.29\pm 0.28}$ $\mathbf{1.18}$ $\mathbf{0.68\pm 0.08}$ $\mathbf{0.63\pm 0.03}$ $\mathbf{0.70\pm 0.09}$ $\mathbf{0.72\pm 0.06}$ $\mathbf{0.68}$ $33.68\pm 14.30$
Figure 3: Visualization of the registration results on one example CT-MR pair in the AbdominalMRCT dataset (posterior-anterior view). Another viewpoint showing the left-to-right perspective is shown in the supplemental material Fig. S2.
HippocampusMR: HD95 and MSD results are reported in Table 3, and the qualitative results are shown in Fig. 4. This dataset includes only the head and tail of the hippocampus, which are relatively small structures with limited deformation compared to abdominal organs. While other non-rigid methods also achieve sub-millimeter MSD, and nr-ICP also achieves sub-millimeter HD95, our method SemICP yields the lowest distance errors. AMM-NRR and Fast-RNRR completed registration on 58 pairs, with two pairs excluded due to numerical issues while building the deformation graph and sampling the influencing node sets.
HD95 MSD Runtime
Head Tail Average Head Tail Average
Initial $3.68\pm 1.46$ $\pm 1.23 3.683.68$ $1.73\pm 0.73$ $1.60\pm 0.48$ $1.67$
ICP $2.40\pm 0.96$ $2.19\pm 0.91$ $2.30$ $1.04\pm 0.35$ $3\pm 0.28 0.990$.99 $\mathbf{0.03\pm 0.00}$
GO-ICP $46\pm 1.04 2.2$$23\pm 1.12$ $2.32$ $1.03\pm 0.37$ $0.94\pm 0.32$ $0.98$ $7.60\pm 2.74$
FR-ICP $2.84\pm 1.43$ $\pm 1.29 2.642.64$ $1.09\pm 0.47$ $0.91\pm 0.31$ $1.00$ $3.46\pm 1.01$
CPD $17\pm 0.78 2.2$$22\pm 0.97$ $2.19$ $0.94\pm 0.24$ $0.92\pm 0.26$ $0.93$ $3.40\pm 0.41$
DCPv1 $10.33\pm 6.22$ $10.99\pm 5.34$ $10.66$ $5.36\pm 4.14$ $4.69\pm 2.48$ $5.03$ $0.07\pm 0.25$
DCPv2 $12.60\pm 6.02$ $13.47\pm 5.41$ $13.03$ $6.81\pm 4.16$ $5.88\pm 2.67$ $6.35$ $0.08\pm 0.26$
nr-ICP $0.77\pm 0.37$ $0.79\pm 0.56$ $0.78$ $0.38\pm 0.11$ $\pm 0.13 0.380.38$ $9.97\pm 2.70$
AMM-NRR $2.61\pm 1.60$ $2.12\pm 1.45$ $2.36$ $0.95\pm 0.50$ $0.76\pm 0.31$ $0.85$ $10.73\pm 2.51$
Fast-RNRR $1.57\pm 1.26$ $1.29\pm 1.13$ $1.43$ $0.62\pm 0.38$ $0.52\pm 0.24$ $0.57$ $6.93\pm 1.31$
SemICP $\mathbf{0.36\pm 0.03}$ $\mathbf{0.36\pm 0.02}$ $\mathbf{0.36}$ $\mathbf{0.17\pm 0.01}$ $\mathbf{0.17\pm 0.01}$ $\mathbf{0.17}$ $4.65\pm 0.61$
Figure 4: Visualization of the registration results on one example MR-MR pair in HippocampusMR dataset (superior-inferior view). Another viewpoint showing the left-to-right perspective is shown in the supplemental material Fig. S3.
3.5 Integration with Learned Labelers
@@
We use nnUNet (Isensee et al. (2021)) as the semantic labeler, since it achieves compatible or superior performance with other state-of-the-art semantic segmentation models on different datasets (Isensee et al. (2024)). After trained nnUNet segments the MR and US volumes, we use the segmented point clouds to register the US point cloud to the MR point cloud using semICP. The goal of the experiment is to show that with easy-to-use segmentation models, our method can register the segmented point clouds that are not as perfect as the ground truth point clouds, instead of focusing on improving the point cloud segmentation quality. More advanced segmentation models may achieve better point cloud segmentation and further improve the registration. Moreover, we use the estimated deformation to warp the US point clouds labeled by an experienced researcher (ground truth point clouds), and evaluate the surface distance between the deformed ground truth US point clouds and the ground truth MR point clouds. This experiment demonstrates that, with our novel deformation representation, the estimated deformation can be readily transferred to new point cloud pairs, a capability that conventional point-matching–based registration methods do not provide. The experiment workflow is shown in Fig. 5.
Figure 5: The proof-of-concept integration with learned labels study.
Table 4: nnUnet learning based segmentation 5-fold evaluation results (Seg vs GT) with Dice and Iou. The metrics are reported in mean (std). MR results are reported on the top two rows, followed by US results at the bottom.
Dataset Metrics Gallbladder IVC Right Kidney Vessels Average
MR T1-W Dice 0.867(0.073) 0.773(0.095) 0.955(0.010) 0.729(0.085) 0.831(0.066)
IoU 0.762(0.088) 0.630(0.106) 0.918(0.012) 0.578(0.096) 0.722(0.075)
US Bmode Dice 0.771(0.069) 0.700(0.076) 0.606(0.091) 0.689(0.073) 0.692(0.072)
IoU 0.670(0.075) 0.560(0.081) 0.520(0.102) 0.543(0.084) 0.573(0.082)
The results of nnUNet segmentation are shown in Table 4, showing that the segmented point clouds are not perfectly accurate. The Dice coefficient is higher for MRI images but lower for US due to the lower image quality and higher noise-to-signal ratio in US. Using the segmented point clouds for registration, the corresponding results are shown in Table 5, and the qualitative example for one MR-US pair is shown in Fig. 6. We first observe that SemICP improves the alignment of segmented point clouds, even though they are not perfectly labeled. Moreover, the results show that the deformation estimated through learning-based labeling and registration also enhances the alignment of the ground-truth point clouds. These findings demonstrate that our proposed method can effectively handle imperfect segmentations and still achieve accurate MR–US alignment.
Table 5: The average HD95 on the liver MR-3D US registration. The metrics are reported in mean (std) in mmmm. Seg refers to the evaluation of the point clouds labeled by learning-based models, and GT refers to the results on the ground truth point clouds.
Initial SemICP-rigid SemICP-final
HD95 MSD HD95 MSD HD95 MSD
Gallbladder Seg $11.21\pm 12.15$ $5.20\pm 8.92$ $8.92\pm 7.21$ $3.96\pm 4.46$ $1.07\pm 0.12$ $0.58\pm 0.06$
GT $9.42\pm 3.89$ $4.19\pm 1.90$ $8.86\pm 4.49$ $3.99\pm 2.68$ $5.06\pm 3.06$ $2.21\pm 1.77$
IVC Seg $11.07\pm 6.60$ $4.54\pm 2.73$ $9.77\pm 8.15$ $4.16\pm 3.63$ $1.36\pm 1.17$ $0.65\pm 0.20$
GT $11.39\pm 7.42$ $4.78\pm 3.41$ $11.48\pm 10.52$ $4.77\pm 4.30$ $7.87\pm 7.54$ $2.85\pm 2.39$
Kidney Seg $21.97\pm 25.26$ $12.14\pm 24.16$ $16.60\pm 12.94$ $7.38\pm 11.86$ $1.17\pm 0.37$ $0.61\pm 0.14$
GT $15.63\pm 5.17$ $7.13\pm 4.71$ $13.99\pm 3.23$ $5.32\pm 2.24$ $6.41\pm 4.62$ $2.49\pm 1.85$
Liver vessel Seg $7.82\pm 5.00$ $2.96\pm 1.53$ $9.16\pm 6.98$ $3.68\pm 2.93$ $1.58\pm 0.37$ $0.74\pm 0.10$
GT $9.92\pm 4.97$ $3.52\pm 1.58$ $11.51\pm 7.17$ $4.37\pm 3.09$ $8.61\pm 5.44$ $2.59\pm 1.53$
Time $7.64\pm 1.12$ $14.37\pm 5.28$
Figure 6: Visualization of the liver case study integrating with learned labelers (anterior-posterior view). Another viewpoint showing the left-to-right perspective is shown in the supplemental material Fig. S4.